OpenStudy (anonymous):

∫_(-2)^3〖m^3 〖 (5+ m^4)〗^7 dm〗evaluate

OpenStudy (mathteacher1729):

YOu can use a u-sub here. :)

OpenStudy (anonymous):

\[\int\limits_{\neg2}^{3}\]

OpenStudy (anonymous):

-2 is the sub

OpenStudy (anonymous):

\[〖m^3 〖 (5+ m^4)〗^7 dm〗\]

OpenStudy (anonymous):

wolfram spits out huge answer for integral

OpenStudy (mathteacher1729):

No I mean "You can use the technique of substitution". let u = 5 + m^4 and then du = 4m^2 dm

OpenStudy (anonymous):

how would you work that out

OpenStudy (anonymous):

Okay. let u=5+m^4 \[\frac{d}{dm}(u=5+m^4)\rightarrow \frac{du}{dm}=5m^3 \implies du=5m^3 dm \implies frac{du}{5}=m^3dm\] Well you have an m^3 dm and a 5+m^4. Replace that making your integral: \[\frac{1}{5} \int\limits u^7 du=\frac{1}{40}u^8 \rightarrow \frac{1}{40}(5+m^4)^8\] Now evaluate it: \[\frac{1}{40}(5+m^4)^8|_{-2}^{3}=\frac{1}{40}[(5+(3)^4)^8-(5+(-2)^4)^8]=9.35*10^{13}\]

OpenStudy (anonymous):

That last part of the first one should read: \[\frac{du}{5}=m^3 dm\]